Mark's soccer team played 12 games. The ratio of games wont to games lost was 2:1. To the nearest whole percent, what percent of the games did his team lose?
Answer by checkley77(12569)

A RATIO OF 2:1 REPRESENTS THE OUTCOME OF THREE EVENTS WHERE 2 WERE WON.
NOW SET UP A RATION WITH FRACTIONS TO REPRESENT THE OUTCOME OF 12 GAMES.
2/3=X/12 NOW CROSS MULTIPLY
3X=2*12
3X=24
X=8 IS THE NUMBER OF WINS OUT OF 12 GAMES PLAYED.
PROOF:
2/1=8/4
8=8